OpenStudy (anonymous):

Mr.Miller received a 10% discount for paying cash for his new car, which was listed for $4695. How much did he actually pay for the car ?

OpenStudy (anonymous):

4695 - 469.5

OpenStudy (anonymous):

$4 225.5

OpenStudy (anonymous):

$4965*0.90 = $4468.50 we use .90 because that is the decimal representation of 90% which would be 100% or the total price minus his 10% discount. another way would be to multiply the original price by 0.10 or the decimal value of 10% to find what ten percent of the original cost would be then subtract that number as shown above by ankur504 so $4965*0.10= 469.5 then subtract that from the original price of $4965-469.5 = $4468.5